A compound of a mineral was found to have the percent composition 29.72% O, 24.65% Fe, and 45.63% Cr. Assuming a 100.0 g sample of the mineral, how many grams are made up of oxygen, iron, and chromium, respectively?


1
Expert's answer
2021-11-22T17:58:02-0500

Assume m = 100 g of sample

Then;

Mass of O = x-O2/100%×mass = 29.47%/100%×100 g = 29.47 g


mass of Fe = x-Fe/100%×mass = 24.70%/100%×100 g = 24.70g


mass of Cr= x-Cr/100%×mass = 45.83 %/100%×100 g = 45.83 g